OpenStudy (calculusxy):

The drawing of a wheel in a book is done at 1/16 scale. If the drawing is 1.8 inches in diameter, what is the wheel's diameter?

OpenStudy (anonymous):

Can you set up a proportion given that information?

OpenStudy (calculusxy):

I don't know how to set up a proportion with this information. If you can guide me, I will try to learn.

OpenStudy (anonymous):

^_^ You'll get it right quick! So, if the wheel in the book is at a scale of 1/16, then we know that every 1 inch of drawing is equivalent to 16 inches of wheel in real life. So then using what we know in the problem we can say \[ \frac{1 \text{ inches in the drawing}}{16 \text{ inches in real life}} = \frac{1.8 \text{ inches in the drawing}}{x \text{ inches in real life}} \] \[\frac{1}{16}=\frac{1.8}{x}\] You just have to solve for x to find the diameter of the real life wheel. Does that make sense?

OpenStudy (calculusxy):

Thank you sooooo much! I have asked people before, but they never gave me an answer like yours. Your awesome. Can you stay with you? I need help with other math ones :)

OpenStudy (anonymous):

yup yup ^^

OpenStudy (calculusxy):

Wait I will mention you in another post.

OpenStudy (anonymous):

k